II. Gains from Trade: Fill the red-colored cells Quansity of Computers Produced 11-1. The graph aside is a production pessibilities fromtier in an economy 1.1 Which point is an inefficient outcome among A, B, C. D, and E? 1,000 2 Which point is an infeasible outcome, which can not be attainabl among A. B, C.D. and E ,200 2000 1-2. Production possibilities frontier shows one trade-off It also shows another of the ten principles, so called opportunity cost 21 When we move from the point B to A. hodution 1,000 computer how many units of cars do we lose 00 600 7001000 Quantityof Cars Produced 22 How many cars are to be sacrificed to get one more unit of e which is asking the opportunity cost of computer at the point B in units of car What is the opportunity cost of computer at the point B in units of car? 1-3. The farmer and rancher each work 8 hours/ day and can devote this time to growing potatoes, rasing cattle, or a combination The farmer can produce an ounce of potatoes in 30m and an ounée of meat in 60 minutes potatoes Farmer The rancher can produce an ounce of potatoes in 10 minutes
Answer #1

Answer:

1.1

Point D is inefficient point because it lies below production possibilities frontier line.

1.2

Point C is infeasible outcome which cannot be attained with current level of resources.

2.1

When we move from B to A we gain 2200-2000=200 units of computers and we lose 700-600=100 units of cars

2.2

Opportunity cost for computer in terms of car =number of cars we lose/ number of computer we gain =100/200=0.5 units of car